# Inferring Dynamic Networks from Marginals with Iterative Proportional Fitting

### v1

@@ 4.1 Structure-dependent MLE Error Bounds
Intuitively, the more “well-connected” a biproportional Poisson network, the better quality its MLEs. We quantify network connectivity through the the Fiedler eigenvalue (Fiedler, 1973), which is the second-smallest eigenvalue $\lambda_{-2}(\mathcal{L})$ of the graph Laplacian $\mathcal{L}:=\mathcal{D}(A\mathbf{1})-A$, where $A$ is the (weighted) adjacency matrix and $\mathcal{D}(\cdot)$ denotes the diagonalization of a vector. In this paper, $\mathcal{L}$ is the graph Laplacian of the weighted bipartite graph $G_{b}$ induced by $\bar{X}$ with $A:=\begin{bmatrix}0&{\bar{X}}\\ {\bar{X}}^{T}&0\end{bmatrix}$. The Fiedler eigenvalue is frequently used in graph theory and distributed optimization to measure graph connectivity (Spielman, 2012). In recent years, its importance for the algorithmic and statistical efficiencies of Luce choice model estimation vis-à-vis the topology of comparison structures has been extensively studied (Shah et al., 2015; Vojnovic & Yun, 2016; Seshadri et al., 2020; Vojnovic et al., 2020; Hendrickx et al., 2020; Bong & Rinaldo, 2022). It is important to note that the Fiedler eigenvalue used in our paper is based on a different graph than in the choice literature, which are constructed from choice data and are not bipartite. An exception is Qu et al. (2023), who provided linear convergence analyses of IPF quantified by the same $\lambda_{-2}(\mathcal{L})$ as in this paper.
@@
which matches the observed deterioration of estimation quality in Figure E.3 as sparsity $r$ increases from 0 to 1.
@@ 5 Guaranteeing IPF Convergence
IPF non-convergence tends to occur when there are many zeros in the inputs (Wong, 1992), and such sparsity is very common in real-world network data, partially due to missing values. For example, in mobility data, true visits may be missed due to noisy GPS signals, populations not carrying cell phones (Coston et al., 2021), or data “clipping” where low values are replaced with zeros to preserve privacy (SafeGraph, 2020a). We discuss these mechanisms in depth in Appendix E.2. Thus, we approach IPF non-convergence from the perspective of missing data: specifically, in initial matrix $X$, and we resolve non-convergence by adding edges to $X$. This is a similar view to the typical solution for IPF non-convergence, which is to replace all zeros in $X$ with very small amounts (Lomax & Norman, 2015). However, the typical approach may result in unrealistic positive entries, such as drivers under the age of 16 when inferring joint demographics (Lovelace et al., 2015) or nonexistent routes in transportation networks. Furthermore, in our setting, replacing all zeros with positive entries completely alters the sparsity structure of the inferred network, which can greatly affect downstream results, such as modeling epidemics (Wang et al., 2003). Thus, we introduce a new algorithm, ConvIPF, that guarantees IPF convergence by adding edges while minimizing changes to the network structure in a specific sense. Even though we focus on network inference, ConvIPF can be used in any application of IPF where non-convergence may be caused by missing values in the initial matrix $X$.
Overview of ConvIPF. Two useful, equivalent conditions that define when IPF converges are (Pukelsheim, 2014):
@@
BLOCKING-SET: identifying a blocking set. Given inputs $X$, $p$, and $q$, where we know IPF does not converge, this subroutine identifies a blocking set of rows $S$ for which Condition (2) is violated. The naive approach to iterate through all subsets until a violation is found, but this approach is extremely inefficient, as there are $2^{m}$ possible subsets. Instead, our subroutine imports ideas from constricted sets in bipartite matching (Hall, 1935; Easley & Kleinberg, 2010) to design a much more efficient algorithm. First, construct a bipartite graph $B$ where the nodes are the rows and columns of $X$ and they are connected wherever $X_{ij}>0$. From running MAX-FLOW to test for convergence, we have flow values for each row/column. Since IPF does not converge, there must be at least one row $i$ whose flow is less than its capacity, $p_{i}$. Run the following variant of breadth-first search (BFS) on $B$ starting from node $n_{i}$. When progressing from a column node $n_{C}$ to its neighboring row node $n_{R}$, only include row nodes where $n_{R}\rightarrow n_{C}$ has non-zero flow in MAX-FLOW. When progressing from row nodes to column nodes, include all (unvisited) neighbors. When BFS terminates, the set of row nodes visited forms a blocking set, which we prove in Appendix D.2.
MODIFY-X: unblocking a blocking set. Given a blocking set $S$, this subroutine minimally adds edges to $X$ to unblock $S$. Let $X^{K}$ represent the modified $X$ after adding new edges $K$ and let $f(X,X^{K})$ represent the change in $X$ that we are trying to minimize. Then, our goal is to find $K^{*}$ that minimizes $f(X,X^{K^{*}})$, subject to $S$ being unblocked under $X^{K^{*}}$, i.e., $\sum_{i\in S}p_{i}\leq\sum_{j\in N_{X^{K^{*}}}(S)}q_{j}$. We consider two natural definitions of $f(X,X^{K})$.
@@ 6 Experiments with Data
Figure 1: Comparing inferred parameters from IPF (x-axis) against inferred parameters from Poisson regression (left y-axis, blue) and true parameters from Poisson model (right y-axis, orange). Grey bars indicate 95% CIs from Poisson regression.
Testing IPF with synthetic data. In our first set of experiments, we use synthetic data generated from our network model (4) to confirm the correspondence between IPF and Poisson regression. In Figure 1, we show that the parameters inferred by IPF and Poisson regression align perfectly, validating Theorem 3.1. Furthermore, our Poisson model enables us to quantify uncertainty in the parameter estimates under the model, adding valuable interpretability missing from IPF output alone. We display 95% confidence intervals (CIs) in Figure 1; while these CIs are only asymptotically valid, they provide useful measures of uncertainty. We also test IPF in more difficult settings, such as with increased sparsity in $\bar{X}$. As expected, we find that the widths of the CIs increase with greater sparsity (Figure 1b), since greater sparsity results in fewer Poisson observations to fit our model (4). We also find that, despite matching the target marginals in all cases, the $\ell_{2}$ distance between IPF’s inferred parameters and the true parameters increases quickly as we increase the sparsity in $\bar{X}$ (Figure E.3). These findings align with Theorem 4.1, where we showed that the bound on the MLEs’ expected estimation error (7) improves as $\bar{X}$ becomes more well-connected.
@@
IPF convergence on mobility data. To test IPF convergence, we use mobility data from SafeGraph (SafeGraph, 2020a). Here, we seek to infer the hourly visit network from neighborhoods to points-of-interest (POIs), so the marginals represent hourly total visits from neighborhoods and to POIs, and $\bar{X}$ represents the time-aggregated network. SafeGraph data provides a real-world example where only the hourly marginals and time-aggregated network are provided, with missing data in the time-aggregated network due to underreporting and data clipping (see Appendix E.2). We use mobility data from the Richmond metro area in Virginia, which has 9917 POIs and 1098 neighborhoods (Chang et al., 2021b). Despite aggregating over 10 months (January to October 2020), $\bar{X}$ remains sparse: only 8% of its entries are non-zero. From running IPF on two days (48 hours), we find that IPF does not converge for three nighttime hours when POI marginals are particularly sparse.
@@
Ground-truth networks from bikeshare data. Using data from New York City’s Citibike system, we can construct ground-truth hourly networks that record the number of bike trips between stations. Acquiring ground-truth networks enables us to test IPF’s ability to infer these networks, compared to baselines, which assesses the downstream utility of our work in bridging IPF to this network inference problem. We test IPF’s inferred networks, given hourly marginals and time-aggregated networks at the month-, week-, and day-level. As a baseline, we try the classic gravity model, which assumes that the amount of travel between two regions is proportional to their inverse distance (Zipf, 1946; Erlander & Stewart, 1990). We fit the “doubly constrained gravity model” (Navick & Furth, 1994), which is equivalent to running IPF on a distance matrix, instead of the time-aggregated network, and the hourly marginals (Appendix E.3.1). This baseline enables us to test how much additional information the time-aggregated network provides, beyond what is captured by common geographical assumptions. We also test other baselines:
@@ 7 Conclusion
In this work, we have established a statistical framework for using IPF to infer a dynamic network from its marginals. Our primary contribution is the biproportional Poisson model. We show that IPF uniquely recovers the MLEs of this model and derive statistical results on the MLEs. Our model not only provides justification for using IPF in this network setting, but also enables new analyses of IPF and clarifies implicit network assumptions. We also introduce ConvIPF, a principled algorithm that guarantees IPF convergence on sparse data, which can be broadly applied wherever IPF is used. Our empirics confirm our theoretical results, demonstrate the value of our modeling and algorithmic contributions, and reveal IPF’s ability to infer networks in practice. Given the long history of IPF and richness of the network inference problem, connecting them generates many avenues for future research. Some future directions include characterizing how IPF’s ability to infer networks varies based on the network’s temporal dynamics (e.g., periodicity), designing variants of IPF that outperform traditional IPF on this network inference problem, and providing a statistical interpretation of our convergence algorithm. Lastly, we note that the biproportional Poisson model is applicable more generally to other instances of the matrix balancing problem beyond network settings.
Broader Impacts
@@
Acknowledgements
S.C. was supported in part by an NSF Graduate Research Fellowship, the Meta PhD Fellowship, and NSF award CCF-1918940. F.K. was supported in part by NSF award CCF-1704417, NSF award IIS1908774, and N. Anari’s Sloan Research Fellowship. Z.Q. was supported in part by ONR Grant N00014-19-1-2468 and NSF CAREER Award IIS-2143176. J.U. was supported in part by NSF CAREER Award IIS-2143176. J.L. was supported in part by NSF awards OAC-1835598, CCF-1918940, DMS-2327709, and Stanford Data Applications Initiative. The authors thank Alfred Galichon and Emma Pierson for helpful discussions and comments as well as Devin Caughey, Guido Imbens, Pang Wei Koh, and members of Jure Leskovec’s lab for helpful feedback on early versions of this work.
@@ Appendix A Details on IPF Implementation
Recall that in the connection between IPF and our network model (4), the number of nonzero entries in $\bar{X}$ is our number of Poisson observations. One implication of this equivalence between modified IPF for non-negative marginals and original IPF on the submatrix is that zero marginals substantially reduce our number of observations, since the submatrix drops entire rows and columns. Thus, for a given hour $t$, our set of observations consists of $\{(i,j)|i\in[m],j\in[n],\bar{X}_{ij}>0,p^{(t)}_{i}>0,q^{(t)}_{j}>0\}$, which can be much fewer than $mn$ observations, given high levels of sparsity in real-world $\bar{X}$, $p^{(t)}$, and $q^{(t)}$. Fewer observations result in less accurate and more uncertain parameter estimates, making explicit the connection between data sparsity and quality of IPF estimates, which we also demonstrate empirically in Section E, especially Figure E.3.
@@ B.1 Duality result for KL divergence minimization
B.2 Proof of Theorem 3.1
@@
B.3 The “joint” network inference problem

### v2

@@ 4.1 Structure-dependent MLE Error Bounds
Intuitively, the more “well-connected” a biproportional Poisson network, the better quality its MLEs. We quantify network connectivity through the Fiedler eigenvalue (Fiedler, 1973), which is the second-smallest eigenvalue $\lambda_{-2}(\mathcal{L})$ of the graph Laplacian $\mathcal{L}:=\mathcal{D}(A\mathbf{1})-A$, where $A$ is the (weighted) adjacency matrix and $\mathcal{D}(\cdot)$ denotes the diagonalization of a vector. In this paper, $\mathcal{L}$ is the graph Laplacian of the weighted bipartite graph $G_{b}$ induced by $\bar{X}$ with $A:=\begin{bmatrix}0&{\bar{X}}\\ {\bar{X}}^{T}&0\end{bmatrix}$. The Fiedler eigenvalue is frequently used in graph theory and distributed optimization to measure graph connectivity (Spielman, 2012). In recent years, its importance for the algorithmic and statistical efficiencies of Luce choice model estimation vis-à-vis the topology of comparison structures has been extensively studied (Shah et al., 2015; Vojnovic & Yun, 2016; Seshadri et al., 2020; Vojnovic et al., 2020; Hendrickx et al., 2020; Bong & Rinaldo, 2022). It is important to note that the Fiedler eigenvalue used in our paper is based on a different graph than in the choice literature, which are constructed from choice data and are not bipartite. An exception is Qu et al. (2023), who provided linear convergence analyses of IPF quantified by the same $\lambda_{-2}(\mathcal{L})$ as in this paper.
@@
which matches the observed deterioration of estimation quality in Figure 2 as sparsity $r$ increases from 0 to 1.
@@ 5 Guaranteeing IPF Convergence
IPF non-convergence tends to occur when there are many zeros in the inputs (Wong, 1992), and such sparsity is very common in real-world network data, partially due to missing values. In fact, we show that, when trying to infer $X^{(t)}$ from $p^{(t)}$, $q^{(t)}$, and $\bar{X}$, for some aggregated time period that includes $t$, IPF will not converge only if there are missing entries in the inputs (Corollary D.1). For example, in mobility data, true visits may be missed due to noisy GPS signals, populations not carrying cell phones (Coston et al., 2021), or data “clipping” where low values are replaced with zeros to preserve privacy (SafeGraph, 2020a). We discuss these mechanisms in depth in Appendix E.2.
Thus, we approach IPF non-convergence from the perspective of missing data: specifically, in initial matrix $X$, and we resolve non-convergence by adding edges to $X$. This is a similar view to the typical solution for IPF non-convergence, which is to replace all zeros in $X$ with very small amounts (Lomax & Norman, 2015). However, the typical approach may result in unrealistic positive entries, such as drivers under the age of 16 when inferring joint demographics (Lovelace et al., 2015) or nonexistent routes in transportation networks. Furthermore, in our setting, replacing all zeros with positive entries completely alters the sparsity structure of the inferred network, which can greatly affect downstream results, such as modeling epidemics (Wang et al., 2003). Instead, we introduce a new algorithm, ConvIPF, that guarantees IPF convergence by adding edges while minimizing changes to the network structure, where we explore two different definitions of network change below. Even though we focus on network inference, ConvIPF can be used in any application of IPF where non-convergence may be caused by missing values in the initial matrix $X$.
Overview of ConvIPF.
Two equivalent conditions that define when IPF converges are (Pukelsheim, 2014):
@@
BLOCKING-SET: identifying a blocking set.
Given inputs $X$, $p$, and $q$, where we know IPF does not converge, this subroutine identifies a blocking set of rows $S$ for which Condition (2) is violated. The naive approach to iterate through all subsets until a violation is found, but this approach is extremely inefficient, as there are $2^{m}$ possible subsets. Instead, our subroutine imports ideas from constricted sets in bipartite matching (Hall, 1935; Easley & Kleinberg, 2010) to design a much more efficient algorithm. First, construct a bipartite graph $B$ where the nodes are the rows and columns of $X$ and they are connected wherever $X_{ij}>0$. From running MAX-FLOW to test for convergence, we have flow values for each row/column. Since IPF does not converge, there must be at least one row $i$ whose flow is less than its capacity, $p_{i}$. Run the following variant of breadth-first search (BFS) on $B$ starting from node $n_{i}$. When progressing from a column node $n_{C}$ to its neighboring row node $n_{R}$, only include row nodes where $n_{R}\rightarrow n_{C}$ has non-zero flow in MAX-FLOW. When progressing from row nodes to column nodes, include all (unvisited) neighbors. When BFS terminates, the set of row nodes visited forms a blocking set, which we prove in Appendix D.2.
MODIFY-X: unblocking a blocking set.
Given a blocking set $S$, this subroutine minimally adds edges to $X$ to unblock $S$. Let $X^{K}$ represent the modified $X$ after adding new edges $K$ and let $f(X,X^{K})$ represent the change in $X$ that we are trying to minimize. Then, our goal is to find $K^{*}$ that minimizes $f(X,X^{K^{*}})$, subject to $S$ being unblocked under $X^{K^{*}}$, i.e., $\sum_{i\in S}p_{i}\leq\sum_{j\in N_{X^{K^{*}}}(S)}q_{j}$. We consider two natural definitions of $f(X,X^{K})$.
@@ 6 Experiments with Data
Figure 1: Comparing inferred parameters from IPF (x-axis) against inferred parameters from Poisson regression (left y-axis, blue) and true parameters from Poisson model (right y-axis, orange). Grey bars indicate 95% CIs from Poisson regression. $N$ represents the number of nonzero entries in $X$, so $N$ is halved with 50% sparsity. Under both networks, estimated parameters from IPF and Poisson regression are perfectly aligned (Theorem 3.1), but their estimation quality worsens with greater sparsity (Theorem 4.1).
Testing IPF with synthetic data.
In our first set of experiments, we use synthetic data generated from our network model (4) to confirm the correspondence between IPF and Poisson regression. In Figure 1, we show that the parameters inferred by IPF and Poisson regression align perfectly, validating Theorem 3.1. Furthermore, our Poisson model enables us to quantify uncertainty in the parameter estimates under the model, adding valuable interpretability missing from IPF output alone. We display 95% confidence intervals (CIs) in Figure 1; while these CIs are only asymptotically valid, they provide useful measures of uncertainty.
Figure 2: Comparing sparsity rate in $\bar{X}$ to number of IPF iterations (left), bound on MLE’s expected estimation error, without constants (middle), and observed $\ell_{2}$ error of IPF estimates (right). Lines represent mean and shaded region represents 95% CIs over 1000 trials.
We also test IPF in more difficult settings, such as with increased sparsity in $\bar{X}$. We generate $\bar{X}$ with a given sparsity rate $r$ by randomly selecting $r\cdot mn$ entries and setting them to 0, and test IPF on varying sparsity rates. As expected, we find that the widths of the CIs increase with greater sparsity (Figure 1b), since greater sparsity results in fewer Poisson observations to fit our model (4). We also find that, despite matching the target marginals in all cases, the $\ell_{2}$ distance between IPF’s inferred parameters and the true parameters increases quickly as we increase the sparsity in $\bar{X}$ (Figure 2, right). These findings align with Theorem 4.1, where we showed that the bound on the MLEs’ expected estimation error (7) improves as $\bar{X}$ becomes more well-connected.
@@
IPF convergence on mobility data.
To test IPF convergence, we use mobility data from SafeGraph (SafeGraph, 2020a). Here, we seek to infer the hourly visit network from neighborhoods to points-of-interest (POIs), so the marginals represent hourly total visits from neighborhoods and to POIs, and $\bar{X}$ represents the time-aggregated network. SafeGraph data provides a real-world example where only the hourly marginals and time-aggregated network are provided, with missing data in the time-aggregated network due to underreporting and data clipping (see Appendix E.2). We use mobility data from the Richmond metro area in Virginia, which has 9917 POIs and 1098 neighborhoods (Chang et al., 2021b). Despite aggregating over 10 months (January to October 2020), $\bar{X}$ remains sparse: only 8% of its entries are non-zero. From running IPF on two days (48 hours), we find that IPF does not converge for three nighttime hours when POI marginals are particularly sparse.
@@
Ground-truth networks from bikeshare data.
Using data from New York City’s Citibike system, we can construct ground-truth hourly networks that record the number of bike trips between stations. Acquiring ground-truth networks enables us to test IPF’s ability to infer these networks, compared to baselines, which assesses the downstream utility of our work in bridging IPF to this network inference problem. We test IPF’s inferred networks, given hourly marginals and time-aggregated networks at the month-, week-, and day-level. As a baseline, we try the classic gravity model, which assumes that the amount of travel between two regions is proportional to their inverse distance (Zipf, 1946; Erlander & Stewart, 1990). We fit the “doubly constrained gravity model” (Navick & Furth, 1994), which is equivalent to running IPF on a distance matrix, instead of the time-aggregated network, and the hourly marginals (Appendix E.3.1). This baseline enables us to test how much additional information the time-aggregated network provides, beyond what is captured by common geographical assumptions. We also test other baselines:
@@ 7 Conclusion
In this work, we have established a statistical framework for using IPF to infer a dynamic network from its marginals. Our primary contribution is the biproportional Poisson model. We show that IPF uniquely recovers the MLEs of this model and derive statistical results on the MLEs. Our model not only provides justification for using IPF to infer dynamic networks, but also enables new analyses of IPF, clarifies implicit network assumptions, and creates natural paths for testing alternate models and misspecification. We also introduce ConvIPF, a principled algorithm that guarantees IPF convergence on sparse data, which can be broadly applied wherever IPF is used. Our empirics confirm our theoretical results, demonstrate the value of our modeling and algorithmic contributions, and reveal IPF’s ability to infer networks in practice. Given the long history of IPF and richness of the network inference problem, connecting them generates many avenues for future research. Some future directions include characterizing how IPF’s ability to infer networks varies based on the network’s temporal dynamics (e.g., periodicity), designing variants of IPF that outperform traditional IPF on this network inference problem, and providing a statistical interpretation of our convergence algorithm. Lastly, we note that the biproportional Poisson model is applicable more generally to other instances of the matrix balancing problem beyond network settings.
Acknowledgements
S.C. was supported in part by an NSF Graduate Research Fellowship, the Meta PhD Fellowship, and NSF award CCF-1918940. F.K. was supported in part by NSF award CCF-1704417, NSF award IIS1908774, and N. Anari’s Sloan Research Fellowship. Z.Q. was supported in part by ONR Grant N00014-19-1-2468 and NSF CAREER Award IIS-2143176. J.U. was supported in part by NSF CAREER Award IIS-2143176. J.L. was supported in part by NSF awards OAC-1835598, CCF-1918940, DMS-2327709, and Stanford Data Applications Initiative. The authors thank Alfred Galichon, Emma Pierson, and Arjun Seshadri for helpful discussions and comments as well as Devin Caughey, Guido Imbens, Pang Wei Koh, and members of Jure Leskovec’s lab for helpful feedback on early versions of this work.
Impact Statement
@@ Appendix A Details on IPF Implementation
Recall that in the connection between IPF and our network model (4), the number of nonzero entries in $\bar{X}$ is our number of Poisson observations. One implication of this equivalence between modified IPF for non-negative marginals and original IPF on the submatrix is that zero marginals substantially reduce our number of observations, since the submatrix drops entire rows and columns. Thus, for a given hour $t$, our set of observations consists of $\{(i,j)|i\in[m],j\in[n],\bar{X}_{ij}>0,p^{(t)}_{i}>0,q^{(t)}_{j}>0\}$, which can be much fewer than $mn$ observations, given high levels of sparsity in real-world $\bar{X}$, $p^{(t)}$, and $q^{(t)}$. Fewer observations result in less accurate and more uncertain parameter estimates, making explicit the connection between data sparsity and quality of IPF estimates, which we also demonstrate empirically in Section E, especially Figure 2.
@@ B.1 Duality result for KL divergence minimization
B.2 Connections to Entropy Regularized Optimal Transport
In many applications of IPF, the initial matrix $X$ arises from some transportation cost function $C$ of the entropic-regularized optimal transport problem associated with (2). More precisely, with the transformation $X=\exp(-C/\varepsilon)$, the KL minimization problem
$\min_{\hat{Y}\in\mathbb{R}^{n\times m}_{+}}D_{\text{KL}}(\hat{Y}\|X)$
$\hat{Y}\mathbf{1}_{m}=p,\quad\hat{Y}^{T}\mathbf{1}_{n}=q$
is equivalent to the following regularized optimal transport problem
$\min_{\hat{Y}\in\mathbb{R}^{n\times m}_{+}}\langle\hat{Y},C\rangle+$ $\varepsilon D_{\text{KL}}(\hat{Y}\|p\otimes q)$
$\hat{Y}\mathbf{1}_{m}=p,$ $\quad\hat{Y}^{T}\mathbf{1}_{n}=q.$
For example, $C$ could be the distance between the origin and the destination or total travel time. In our setting, $X$ is given as $\bar{X}$, the time-aggregated network (e.g., number of bike trips between stations over time or number of visits from neighborhoods to places), but we could still interpret it as a surrogate measure for the above mentioned physical quantities. For example, we can expect that the longer it takes to travel between a particular pair of locations, the fewer people on average will travel between those nodes. Therefore, when we do not have access to information about distance or travel cost, $\bar{X}$ can be used as a reasonable surrogate.
We also prefer to directly use $\bar{X}$ instead of writing it as $\exp(-C/\epsilon)$ since in real-world data, $\bar{X}$ contains many zero elements. If we want to interpret it in terms of $C$, we need to allow “infinite transportation cost” for some origin-destination pairs. In Section E.1.3, we also briefly discuss a possible extension of our biproportional Poisson model based on the gravity model of Navick & Furth (1994), which uses the distance between origin-destination pairs as the cost function $C$, and with $\bar{X}$ the aggregate number of trips as a constant multiplier instead. When we have access to additional information (besides aggregate trips) that can serve as the cost $C$, such a model could be preferable.
B.3 Proof of Theorem 3.1
@@ B.2 Connections to Entropy Regularized Optimal Transport
Comparison to collective graphical models.
Sheldon & Dietterich (2011) introduce the problem of collective graphical models (CGMs): how to fit a model of individual behavior from aggregate data (e.g., counts). The general problem that CGM addresses shares strong similarities with the dynamic network inference problem in our work, where we are also trying to estimate finer-grained information (time-varying networks) based on coarser, aggregate data (time-varying marginals and time-aggregated network) due to privacy or data collection costs. A common feature is that the available aggregate data provides the sufficient statistics of the model, which enables valid estimation and inference. An additional algorithmic connection between our work and CGM is that IPF has also been used in the inference of CGM (Singh et al., 2020).
However, a notable difference is that CGM is typically applied in settings where only the time-varying marginals are known, without access to the time-aggregated network. For example, several papers use CGMs to estimate population flows given the number of people in each area over time (Iwata et al., 2017; Akagi et al., 2018; Iwata & Shimizu, 2019). Furthermore, CGM seeks to relate population-level observations to a model of individual behavior (e.g., a single person’s mobility or a bird’s migration patterns), while we are not trying to estimate individual models. Instead of disaggregating over populations, we seek to disaggregate over time, by taking the time-aggregated network and distributing it over smaller time intervals (e.g., from monthly to hourly). Finally, the specific mathematical formulations of the problems in CGM and in the network inference problem seem to be distinct. CGM is built on probabilistic graphical models of independent and identically distributed individuals. It aims to estimate the parameters of random vectors on nodes of a graph given only aggregate counts of specific states using MAP inference. In contrast, our biproportional Poisson model is used to model the random population flows between nodes of a graph using maximum likelihood estimation, which is solved via IPF.
B.4 The “joint” network inference problem